If an event occurs 38 times out of 100 samples, the frequency of that event is
Answer #1

Frequency means no. of times an event occurred in total possibilities. For example, if of 100 puppies, only 50 survive to maturity. So frequency of survivors = 50 out of 100 = 50/100= 0.5
So in given question, 38 times event occurred out of 100.

So frequency = 38 in 100. = 38/100 = 0.38

If you want in percentage, then multiply frequency by 100. --> 0.38*100 = 38%